# Heuristic Reasoning in AI: Instrumental Use and Mimetic Absorption

### 14 Mar 2024 v1

@@ Abstract
We propose a novel program of heuristic reasoning within artificial intelligence (AI) systems. Through a series of innovative experiments, including variations of the classic Linda problem and a novel application of the Beauty Contest game, we uncover trade-offs between accuracy maximization and effort reduction that shape the conditions under which AIs transition between exhaustive logical processing and the use of cognitive shortcuts (heuristics). We distinguish between the ‘instrumental’ use of heuristics to match resources with objectives, and ‘mimetic absorption,’ whereby heuristics are learned from humans, and manifest randomly and universally. We provide evidence that AI, despite lacking intrinsic goals or self-awareness, manifests an adaptive balancing of precision and efficiency, consistent with principles of resource-rational human cognition as explicated in classical theories of bounded rationality and dual-process theory.
@@ Introduction
Heuristics in human cognition—cognitive shortcuts that facilitate mental processing—are situated within contrasting narratives. Simon’s notion of bounded rationality (Simon 1955) casts heuristics as tools that enable navigation in environments too complex for the unaided mind. When aligned with psychological capacities and grounded in ecological rationality, a parallel view advocates for a ‘fast and frugal’ approach to cognition (Gigerenzer and Goldstein 1996), where heuristics serve as scaffolds in decisions that might prove unnecessary, intractable, or suboptimal if reliant solely on analytic processing (Simon 1956). In contrast, a ‘heuristics as bias’ view frames heuristics as leading to systematic and predictable deviations from optimal decision-making, given standards of complete information processing (Gilovich et al. 2002, Tversky and Kahneman 1974). Implicit in the latter is the attainability of a decision through complete analytic processing such that the use of a shortcut yields a suboptimal outcome—in the face of situations where analytic processing is infeasible, clearly, a heuristic can yield a better decision than the roll of a dice.
Our paper proposes a novel program of heuristic reasoning as it applies to artificial intelligence (AI) cognition. Contemporary AI is designed without a sense of self and lacks introspective capabilities. It operates within the constraints of computational resources allocated for each iteration of its response during dialogue with a human trainer in reinforcement training, or within the limits of its context window—factors that act as implicit resource controls. Yet, it is capable of engaging in complex reasoning by breaking down problems into interconnected tasks—a process akin to chain-of-thought reasoning—though it lacks the capacity for truly deliberate and effort-intensive thinking, as it does not possess a concept of effort.
@@
We posit that whereas human cognition reflects the use of dual systems due to intrinsic constraints shaped by evolution, machine learning algorithms (including AI) conduct searches for efficacy during training, retaining functions that maximize objectives within their computational resource limits. As human cognition is shaped by nature and nurture, AI cognition is tied to its programming and training. Consequently, AI’s cognitive strategies also evolve in response to objectives and training. For instance, a low-capacity AI trained to play 100-dimensional chess, a game requiring high-capacity strategy, may rely on heuristics, while a high-capacity AI trained to play tic-tac-toe, a game requiring low-capacity strategy, may express only optimal strategies. An AI trained on a multitude of scenarios may employ both elaborate processing and mental shortcuts. Similar to the adaptive strategies observed in human cognition (Evans 2008, Kahneman 2003), it may dynamically optimize decision-making precision by selectively employing or discarding heuristics, actively refining the set of considered solutions to efficiently manage computational load.
@@
Specifically, an AI may have learned during its training to employ heuristics selectively, transitioning to more elaborate processing when the prompt contains information—processing cues—that signal feasibility (i.e., absence of resource constraints) and necessity (i.e., need for precision). We term such use ‘instrumental,’ as a switching process, even when activated reflexively, is designed to optimally match resources with objectives. In contrast, if heuristics are absorbed mimetically from human data and interactions, they may manifest universally. For instance, in an AI trained to play strategy games, an instrumental absorption of heuristics would correspond to the AI learning cues (e.g., whether the game is 100-dimensional chess or tic-tac-toe) that allow it to determine which strategy (heuristic or analytic processing) is likely to be more beneficial. Conversely, a mimetic absorption of heuristics would imply that heuristic processing may emerge universally and randomly, irrespective of which game (100-dimensional chess or tic-tac-toe) is being played and what the AI’s processing resources are.
@@
Our work contributes to an emerging body of literature on AI cognition—the capability of AI to perceive, understand, reason, and learn from information. Prior evidence indicates that while previous generations of AIs (e.g., OpenAI’s GPT-3) underperformed on human psychological assessments, contemporary AI (e.g., GPT-4) demonstrates performance comparable to humans (Trott et al. 2023). Further evidence pertains to causal reasoning, encompassing abstract reasoning (Webb et al. 2023) and inductive reasoning (Han et al. 2024). Nonetheless, other studies suggest that such abilities may be attributed to lexical cues. For example, emphasizing a reliance on rote memorization, Ullman (2023) exposes the vulnerability of AI to even minor shifts in the linguistic structure of established assessments. These limitations echo critiques from Chomsky et al. (2023) and Pearl and Mackenzie (2018), as well as theoretical results by Fodor and Pylyshyn (1988), pointing to fundamental obstacles within connectionist architectures that hinder an AI’s understanding of complex causal explanations and suggest a reliance on the repetition of learned responses.
@@
Critically, current assessments frequently rely on classic psychological tests and assessments tailored for humans, such as the Torrance Tests (Guzik et al. 2023). This reliance introduces several limitations: (1) AIs, being trained on performance benchmarks that encompass psychology and sociology, may already be familiar with the expected outcomes of a test before taking it, complicating the distinction between genuine responses and rote memorization; (2) Assessments may fail to provide conclusive insights due to the inherent challenge of accessing the AI’s introspective processes—specifically, understanding the rationale behind its decisions; (3) Questions related to a sense of self, such as ‘What would you choose?’ pose a significant challenge for contemporary AI, which lacks an innate sense of self and has no utility for making choices; (4) Many assessments are structured around scenarios that relegate the AI to the role of a passive observer of human interactions (Bubeck et al. 2023). As a result, while these studies shed light on the AI’s ability to mimic human Theory of Mind (ToM) traits (Langley et al. 2022) in scenario-based analyses, they fall short of thoroughly examining the AI’s decision-making in contexts that require adaptive responses to varying levels of decision complexity and computational resources.
@@ Discussion
We provide evidence that showcases heuristic use by AI in specific contexts and circumstances. Distinct from prior examinations of AI, we seek to establish authentic responses. Therefore, we innovate by constructing experimental conditions that draw on established techniques but introduce novelties aimed at overcoming the tendency of AI to rely on memorization—a facet of AI’s capabilities that is its known advantage (Bender et al. 2021).
In particular, we distinguish between the mimetic absorption of heuristics and their instrumental utilization. These mechanisms relate to their source: whereas mimetic absorption pertains to the imitation of patterns in human interactions and human-generated data—for instance, by emulating a human interlocutor who demonstrates heuristic use and System 1 processing, the employment of heuristics as an instrument points to intrinsic optimizations whereby environmental regularities shape efficient but biased cognitive processes.
@@
Finally, we situate our conceptualization in AI training. An alternative view may reflect on the fact that neural networks are functional models of biological brains. They represent words and concepts numerically and form responses by varying the attention they pay to different representations (Vaswani et al. 2017). This mechanism bears a striking resemblance to the representativeness heuristic. To the extent that our observations of the representativeness heuristic emerge from the fundamentals of a connectionist architecture, similar cognitive biases may arise naturally in AIs, even if absent in their training. This intersection of AI design, AI cognition, and cognitive patterns offers a promising avenue for future research (see Szegedy et al. 2013 for similar observations in computer vision).

### 18 Mar 2024 v2

@@ Abstract
Deviating from conventional perspectives that frame artificial intelligence (AI) systems solely as logic emulators, we propose a novel program of heuristic reasoning. We distinguish between the ‘instrumental’ use of heuristics to match resources with objectives, and ‘mimetic absorption,’ whereby heuristics manifest randomly and universally. Through a series of innovative experiments, including variations of the classic Linda problem and a novel application of the Beauty Contest game, we uncover trade-offs between maximizing accuracy and reducing effort that shape the conditions under which AIs transition between exhaustive logical processing and the use of cognitive shortcuts (heuristics). We provide evidence that AIs manifest an adaptive balancing of precision and efficiency, consistent with principles of resource-rational human cognition as explicated in classical theories of bounded rationality and dual-process theory. Our findings reveal a nuanced picture of AI cognition, where trade-offs between resources and objectives lead to the emulation of biological systems, especially human cognition, despite AIs being designed without a sense of self and lacking introspective capabilities.
@@ Introduction
Heuristics in human cognition—cognitive shortcuts that facilitate mental processing—are situated within contrasting narratives. Simon’s notion of bounded rationality (Simon 1955) casts heuristics as tools that enable navigation in environments too complex for the unaided mind. When aligned with psychological capacities and grounded in ecological rationality, a parallel view advocates for a ‘fast and frugal’ approach to cognition (Gigerenzer and Goldstein 1996), where heuristics serve as scaffolds in decisions that might prove unnecessary, intractable, or suboptimal if reliant solely on analytic processing (Simon 1956). In contrast, a ‘heuristics as bias’ view frames heuristics as leading to systematic and predictable deviations from optimal decision-making, given standards of complete information processing (Gilovich et al. 2002, Tversky and Kahneman 1974). Implicit in the latter perspective is the assumed feasibility of complete analytic processing—the use of a shortcut only yields a suboptimal outcome (i.e., biased decision-making leads to suboptimal outcomes) if the optimal is achievable; clearly in situations where analytic processing is infeasible, a heuristic can yield a better decision than random chance.
Drawing from human cognition, our paper proposes a novel program of heuristic reasoning as it applies to artificial intelligence (AI) cognition. Given that AIs lack the capacity for truly deliberate and effort-intensive thinking and do not possess a concept of effort, conventional thinking dictates that AI engages in the emulation of exhaustive logic and rational reasoning. To this end, a failure in reasoning is often seen as a sign of weakness in programming, a deficiency in computational abilities, or a lack of information.
@@
In contrast, we posit that whereas human cognition reflects the use of dual systems due to intrinsic constraints shaped by evolution, machine learning algorithms (including AI) conduct searches for efficacy during training, retaining functions that maximize objectives within their computational resource limits. As human cognition is shaped by both nature and nurture, AI cognition is similarly tied to its programming and training.
Consequently, AI’s cognitive strategies evolve in response to objectives and training. For instance, a low-capacity AI trained to play 100-dimensional chess, a game requiring high-capacity strategy, may rely on heuristics, while a high-capacity AI trained to play tic-tac-toe, a game requiring low-capacity strategy, may express only optimal strategies. An AI trained on a multitude of scenarios may employ both elaborate processing and mental shortcuts. Similar to the adaptive strategies observed in human cognition (Evans 2008, Kahneman 2003), it may dynamically optimize decision-making precision by selectively employing or discarding heuristics, actively refining the set of considered solutions to efficiently manage computational load. Specifically, it may have learned during training to employ heuristics selectively, transitioning to more elaborate processing when the prompt contains information—processing cues—that signal feasibility (i.e., absence of resource constraints) and necessity (i.e., need for precision).
@@
We term such use ‘instrumental,’ as a switching process, even when activated reflexively, is designed to optimally match resources with objectives. In contrast, if heuristics are absorbed mimetically from human data and interactions, they may manifest universally. For instance, in an AI trained to play strategy games, instrumental absorption of heuristics would correspond to the AI learning cues (e.g., whether the game is 100-dimensional chess or tic-tac-toe) that allow it to determine which strategy (heuristic or analytic processing) is likely to be more beneficial. Conversely, mimetic absorption of heuristics would imply that heuristic processing may emerge universally and randomly, irrespective of the specific game (100-dimensional chess or tic-tac-toe) being played and the AI’s processing resources.
@@
Our work contributes to an emerging body of literature on AI cognition—the capability of AIs to perceive, understand, reason, and learn from information. Prior evidence indicates that while previous generations of AIs (e.g., OpenAI’s GPT-3) underperformed on human psychological assessments, contemporary AI (e.g., GPT-4) demonstrates performance comparable to humans (Trott et al. 2023). Further evidence pertains to causal reasoning, encompassing abstract reasoning (Webb et al. 2023) and inductive reasoning (Han et al. 2024). Nonetheless, other studies suggest that such abilities may be attributed to lexical cues. For example, emphasizing a reliance on rote memorization, Ullman (2023) exposes the vulnerability of AI to even minor shifts in the linguistic structure of established assessments. These limitations echo critiques from Chomsky et al. (2023) and Pearl and Mackenzie (2018), as well as theoretical results by Fodor and Pylyshyn (1988), pointing to fundamental obstacles within connectionist architectures that hinder an AI’s understanding of complex causal explanations and suggest a reliance on the repetition of learned responses.
@@
Critically, current assessments rely on classic psychological tests and assessments tailored for humans, such as the Torrance Tests (Guzik et al. 2023). This reliance introduces several limitations: (1) AIs, being trained on performance benchmarks that encompass psychology and sociology, may already be familiar with the expected outcomes of a test before taking it, complicating the distinction between genuine responses and rote memorization; (2) Assessments may fail to provide conclusive insights due to the inherent challenge of accessing the AI’s introspective processes—specifically, understanding the rationale behind its decisions; (3) Questions related to a sense of self, such as ‘What would you choose?’ pose a significant challenge for contemporary AI, which lacks an innate sense of self and has no utility for making choices; (4) Many assessments are structured around scenarios that relegate the AI to the role of a passive observer of human interactions (Bubeck et al. 2023). As a result, while these studies shed light on the AI’s ability to mimic human Theory of Mind (ToM) traits (Langley et al. 2022) in scenario-based analyses, they fall short of thoroughly examining the AI’s decision-making in contexts that require adaptive responses to varying levels of decision complexity and computational resources.
@@ Discussion
We provide evidence showcasing heuristic use by AI in specific contexts and circumstances. We seek to establish authentic responses. Therefore, we innovate by constructing experimental conditions that draw on established techniques but introduce novelties aimed at overcoming the tendency of AI to rely on memorization—a facet of AI’s capabilities that is its known advantage (Bender et al. 2021).
We distinguish between the mimetic absorption of heuristics and their instrumental utilization. These mechanisms relate to their source: whereas mimetic absorption pertains to the imitation of patterns in human interactions and human-generated data—for instance, by emulating a human interlocutor who demonstrates heuristic use and System 1 processing, the employment of heuristics as an instrument points to intrinsic optimizations whereby environmental regularities shape efficient but biased cognitive processes.
@@
We situate our conceptualization in AI training. An alternative view may reflect that neural networks are functional models of biological brains. They represent words and concepts numerically and form responses by varying the attention they pay to different representations (Vaswani et al. 2017). This mechanism bears a striking resemblance to the representativeness heuristic. To the extent that our observations of the representativeness heuristic emerge from the fundamentals of a connectionist architecture, similar cognitive biases may arise naturally in AIs, even if absent in their training. This intersection of AI design, AI cognition, and cognitive patterns offers a promising avenue for future research (see Szegedy et al. 2013 for similar observations in computer vision).